The weights of four randomly and independently selected bags of tomatoes labeled 5.0 pounds were found to be 5.4​, 5.0​, 5.3​, and 5.5 pounds. Assume Normality. Answer parts​ (a) and​ (b) below.

a. Find the 95% confidence interval for the mean weight of all bags of tomatoes. (____ , ____)

b. Can you reject the population mean of 5 pounds?
Answer #1

sample mean, xbar = 5.3
sample standard deviation, s = 0.216
sample size, n = 4
degrees of freedom, df = n - 1 = 3

Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 3.182

CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (5.3 - 3.182 * 0.216/sqrt(4) , 5.3 + 3.182 * 0.216/sqrt(4))
CI = (4.96 , 5.64)

b)
can not reject the population mean of 5 pounds